OpenStudy (anonymous):

(2)(4)-(6-4) divided by -3

OpenStudy (rock_mit182):

operate what is inside the parentheses, then multiply 4 and 2, then substract, finally divide by 3

OpenStudy (anonymous):

\[\frac{ (2)(4)-(6-4) }{ -3 }\] \[=\frac{ 8 - 6 + 4 }{ -3 }\] \[= \frac{ 6 }{ -3 }\] \[= - 2\]

OpenStudy (rock_mit182):

you did her homework lol, thats better

OpenStudy (anonymous):

:)